the formula for the volume of a cone is V=1/3Bh. If the Volume, V, is 15 cubic centimeters and the height, h, is 5 centimeters, find the base, B, of this cone.
Answer by ewatrrr(24785)

 
Hi
V = 1/3Bh
15cm^3 = (1/3)B*5cm
45cm^3/5cm = B
9cm^2 = B